# Covariance test and universal bootstrap by operator norm

### v1

@@ 3.2 Universality
We provide several remarks regarding Assumption 4. Informally, condition (27) ensures that the extreme eigenvalues of $\bm{\Sigma}$ do not spread near the endpoints $E_{\pm}$, thereby preventing spikes outside the support of $\rho$. Similar assumptions have been made in the literature for the universality of $\hat{\bm{\Sigma}}$, using the Stieltjes transform in place of $\tilde{m}$ and $\lambda_{i}(\bm{\Sigma})$ in place of $\lambda_{i}(\bm{\Sigma}(E_{\pm}))$, as in Bao, Pan and Zhou (2015); Knowles and Yin (2017). This aligns with the intuition that the effect of $\bm{R}$ can be expressed through the transformation from $\bm{\Sigma}$ to $\bm{\Sigma}(z)$. Moreover, (27) is automatically satisfied when $\phi=p/n\to\infty$, i.e. $\alpha>1$. We thus impose Assumption 4 only in the case $\alpha=1$.
To state our main results, we define the concept of stochastic domination, introduced by Erdős, Knowles and Yau (2013) and widely applied in random matrix theory (Knowles and Yin, 2017). For two families of non-negative random variables $A=\left(A^{(n)}(u)\ :\ n\in\mathbb{N},u\in U^{(n)}\right)$ and $B=\left(B^{(n)}(u)\ :\ n\in\mathbb{N},u\in U^{(n)}\right)$, where $U^{(n)}$ is $n$-dependent parameter set, we say $A$ is stochastically dominated by $B$ uniformly in $u$ if for any $\epsilon>0$ and $D>0$, we have $\sup_{u\in U^{(n)}}\mathbb{P}\left(A^{(n)}(u)>n^{\epsilon}B^{(n)}(u)\right)\leq n^{-D}$ for large enough $n\geq n_{0}(\epsilon,D)$. We use the notation $A\prec B$ to represent this relationship. When $A$ is a family of negative or complex random variables, we also write $A\prec B$ or $A=O_{\prec}(B)$ to indicate $|A|\prec B$. With these definitions, we state the anisotropic local law as follows.
Theorem 3.1 (Anisotropic local law).
Define the control parameter as $\Psi(z)=\sqrt{\frac{\Im\tilde{m}(z)}{n\eta}}+\frac{1}{n\eta}$. Suppose that Assumptions 1, 2, 3, 4 and (17) hold.
(i) We have
$\bm{u}^{T}\bm{\Phi}\bar{\bm{\Sigma}}^{-1}(\bm{G}(z)-\bar{\bm{G}}(z))\bar{\bm{\Sigma}}^{-1}\bm{\Phi}\bm{v}=O_{\prec}(\|\bm{u}\|\|\bm{v}\|\Psi(z)),$ (28)
uniformly in vectors $\bm{u},\bm{v}\in\mathbb{R}^{n+2p}$ and $z\in\mathcal{D}$, where the weight matrix $\bm{\Phi}$ is defined as $\bm{\Phi}=\text{diag}(\phi^{-1/4}\bm{I}_{p},\phi^{-3/4}\bm{I}_{p},\bm{I}_{n})$ and the augemented covariance matrix is $\bar{\bm{\Sigma}}=\text{diag}(\bm{\Sigma},\bm{I}_{n+p})$.
(ii) Moreover, we have the average local law
$m_{p}(z)-\bar{m}_{p}(z)=O_{\prec}((n\eta)^{-1}),$ (29)
uniformly in $z\in\mathcal{D}$.
The anisotropic local law provides a delicate characterization of the discrepancy between the random Green function and its deterministic counterpart, yielding a precise bound to analyze the behavior of the extreme eigenvalues of $\bm{M}_{n}^{\prime}$. When $\bm{R}=\bm{0}$, similar results have been established for $\alpha=1$ in Knowles and Yin (2017) and for $\alpha\geq 1$ in Ding and Wang (2023); Ding, Hu and Wang (2024). As demonstrated in Ding and Wang (2023), the results for general $\alpha\geq 1$ hold without requiring the dimension $p$ and sample size $n$ to grow proportionally, resulting in different convergence rates for the blocks of $\bm{G}(z)$. Ding and Wang (2023) separately provides convergence rates for each block and offers a coarse bound for the anisotropic local law
$G_{i\mu}(z)=O_{\prec}(\phi^{-\frac{1}{4}}\Psi(z)),\ i=1,\cdots,p,\mu=p+1,\cdots,2p+n.$
In this study, we enhance these findings by introducing the weight matrix $\bm{\Phi}$, allowing us to express the anisotropic local law (28) in a more compact form. This reformulation refines the convergence rate and simplifies our proof.
Building on the anisotropic local law theorem, we proceed to establish our universality result. Our first step involves deriving key implications from Theorem 3.1. Recall that $\rho$ is the measure on $\mathbb{R}$ whose Stieltjes transform is the deterministic equivalence $\bar{m}_{p}(z)$. We define the quantile sequence of $\rho$ as $w_{1}\geq\cdots\geq w_{p}$ such that $\int_{w_{i}}^{+\infty}\rho(x)\mathrm{d}x=(i-1/2)/p$ for $i=1,\cdots,p$. We aim to demonstrate that the eigenvalues $\lambda_{i}(\bm{M}_{n}^{\prime})$ of $\bm{M}_{n}^{\prime}$ remain close to $w_{i}$ for $i=1,\cdots,p$.
Theorem 3.2 (Rigidity of eigenvalues).
Suppose that Assumptions 1, 2, 3, 4 hold.
(i) When $\alpha=1$, we have for any fixed integer number $k\geq 1$,
$|\lambda_{i}(\bm{M}_{n}^{\prime})-w_{i}|,\ |\lambda_{p-i}(\bm{M}_{n}^{\prime})-w_{p-i}|\prec(\min\{i,p+1-i\})^{-\frac{1}{3}}n^{-\frac{2}{3}},$ (30)
uniformly for $1\leq i\leq k$.
(ii)When $\alpha>1$, we have
$|\lambda_{i}(\bm{M}_{n}^{\prime})-w_{i}|\prec(\min\{i,n+1-i\})^{-\frac{1}{3}}n^{-\frac{2}{3}},$ (31)
uniformly for $1\leq i\leq n$.
Theorem 3.2 establishes the rigidity of the eigenvalues of $\bm{M}_{n}^{\prime}$, with two main implications. First, when $\alpha=1$, equation (30) shows that the largest and smallest $k$-th eigenvalues lie within an $n^{-2/3+\epsilon}$-neighborhood of $w_{p}$, respectively, for any $\epsilon>0$. Furthermore, we have $|w_{1}-E_{+}|,|w_{p}-E_{-}|\prec n^{-2/3}$, leading to $|\lambda_{i}(\bm{M}_{n}^{\prime})-E_{+}|,|\lambda_{p-i}(\bm{M}_{n}^{\prime})-E_{-}|\prec n^{-2/3}$ uniformly for $1\leq i\leq k$. Second, with Theorem 3.2, we demonstrate that $w_{i}\asymp\phi^{1/2}\to+\infty$ for $1\leq i\leq n$, revealing a fast $n^{-2/3+\epsilon}$ approximation rate for the diverging quantities $\lambda_{i}(\bm{M}_{n}^{\prime})$ and $w_{i}$.
@@
To establish the universality result, we provide bounds for the discrepancy between the distribution of $\lambda_{1}(\bm{M}_{n}^{\prime})$ and its Gaussian counterpart. We denote by $\mathbb{P}^{\text{Gau}}$ and $\mathbb{E}^{\text{Gau}}$ the probability and expectation under the additional assumption that $\{Z_{ij}\}_{1\leq i\leq n,1\leq j\leq p}$ are independent standard Gaussian variables. With this notation, we present the following result.
@@
$\mathbb{P}(n^{\frac{2}{3}}($ $\lambda_{1}(\bm{M}_{n}^{\prime})-E_{+})\leq t-n^{-\epsilon})-n^{-\frac{1}{6}+C\epsilon}$
$\leq$ $\mathbb{P}^{\text{Gau}}(n^{\frac{2}{3}}(\lambda_{1}(\bm{M}_{n}^{\prime})-E_{+})\leq t)$ (32)
$\leq\mathbb{P}(n^{\frac{2}{3}}(\lambda_{1}(\bm{M}_{n}^{\prime})-E_{+})\leq t+n^{-\epsilon})+n^{-\frac{1}{6}+C\epsilon}.$
When $\alpha=1$, similar results also hold for $\lambda_{p}(\bm{M}_{n}^{\prime})$.
The universality Theorem 3.3 shows that the asymptotic distributions of $\lambda_{1}(\bm{M}_{n})$, $\lambda_{p}(\bm{M}_{n})$, and consequently $\sigma_{1}(\bm{M}_{n})=\|\bm{M}_{n}\|_{\text{op}}$, rely solely on the first two moments of $\bm{X}$. In contrast, as discussed in Section 4.3, other widely-used norms of $\bm{M}_{n}$, such as $\|\bm{M}_{n}\|_{\text{F}}$ and $\|\bm{M}_{n}\|_{\text{sup}}$,are influenced by the first four moments of $\bm{X}$. This universality characteristic of the operator norm offers a straightforward yet effective framework for constructing rejection regions in hypothesis testing. A detailed exploration of this application is presented in Section 3.3. Before applying universality to covariance matrix testing, we outline several corollaries of Theorem 3.3, which are of notable independent interest.
@@
Consider the case $\alpha>1$, $\bm{\Sigma}=\bm{I}_{p}$, $\bm{R}=\bm{0}$. Define $\mu=p$, $\sigma=p^{1/2}n^{-1/6}$. Under Assumption 1, we have as $n\to\infty$,
@@
where $\Rightarrow$” represents convergence in law, and $\text{TW}_{1}$ is the Tracy-Widom distribution of type $1$.
One of the central problems in random matrix theory is establishing the asymptotic distribution of the largest eigenvalue of a sample covariance matrix. In this context, (33) was derived for Gaussian $\bm{Z}$ in the setting where $p/n\to\infty$ (Karoui, 2003). However, for general distributions of $\bm{Z}$ where $\alpha>1$, no further results have been established. Our universality Theorem 3.3 addresses this significant gap in random matrix theory.
Corollary 2.
Consider the case $\alpha=1$, $\bm{\Sigma}=\bm{I}_{p}$, and a general symmetric matrix $\bm{R}$. Under Assumptions 1,2,4, we have $n\to\infty$,
$\frac{n^{\frac{2}{3}}(\lambda_{1}(\hat{\bm{\Sigma}}+\bm{R})-E_{+})}{\sigma}\Rightarrow\text{TW}_{1}.$ (34)
Here $E_{+}$ and $\sigma$ are two constants depending on $\bm{R}$.
As demonstrated in Section 5 in the Supplement, when $\bm{Z}$ follows a Gaussian distribution, (34) can be derived using the findings from Ji and Park (2021). For non-Gaussian distributions of $\bm{Z}$, our universality Theorem 3.3 extends this result to the same form. These results expand the traditional analysis of extreme eigenvalues in the sample covariance matrix $\bm{\hat{\Sigma}}$ to a more general setting $\hat{\bm{\Sigma}}-\bm{\Sigma}_{0}$, where $\bm{R}=-\bm{\Sigma}_{0}$, with broader applications in covariance testing within high-dimensional statistics.
3.3 Universal bootstrap for covariance test
Theorem 3.3 is commonly termed universality in random matrix literature concerning extreme eigenvalue, see Bao, Pan and Zhou (2015); Knowles and Yin (2017). However, this universality does not ensure the validity of our universal bootstrap procedure, which depends on bounds like (7) and (8). Specifically, (3.3) provides bounds on $\mathbb{P}(n^{2/3}(T-\phi^{1/2}E)\leq t\pm n^{-\epsilon})$ and $\mathbb{P}(n^{2/3}(T^{\text{ub}}-\phi^{1/2}E)\leq t)$ with $E=\max\{E_{+},-E_{-}\}$, while we aim to bound $\mathbb{P}(n^{2/3}(T-\phi^{1/2}E)\leq t)$ and $\mathbb{P}(n^{2/3}(T^{\text{ub}}-\phi^{1/2}E)\leq t)$. Establishing this bound requires demonstrating the closeness between $\mathbb{P}(n^{2/3}(T-\phi^{1/2}E)\leq t\pm n^{-\epsilon})$ and $\mathbb{P}(n^{2/3}(T-\phi^{1/2}E)\leq t)$. This step relies on the anti-concentration inequality (Chernozhukov, Chetverikov and Kato, 2013) that will be established below.
We first define the Lévy anti-concentration function for random variable $T$ as
$a(T,\delta)=\sup_{t\in\mathbb{R}}\mathbb{P}(t\leq T\leq t+\delta),$ (35)
for $\delta>0$. We shall provide bounds for $a(T,\delta)$. We note that the theorems derived thus far apply specifically to the extreme eigenvalues $\lambda_{1}(\bm{M}_{n}^{\prime})$ and $\lambda_{p}(\bm{M}_{n}^{\prime})$. As we will show, extending these results to extreme singular values $\sigma_{1}(\bm{M}_{n}^{\prime})$ requires only weaker assumptions.
@@ 3.3 Universal bootstrap for covariance test
When $\alpha=1,$ we require that there exists $\tau>0$ such that
@@
Assumption 4 is evidently weaker than Assumption 4. The reasoning behind Assumption 4 is straightforward: given that the singular value $\sigma(\bm{M}_{n}^{\prime})=\max\{|\lambda_{1}(\bm{M}_{n}^{\prime})|,|\lambda_{p}(\bm{M}_{n}^{\prime})|\}$, if $|E_{+}|>|E_{+}|$, then with probability $1$, $\sigma(\bm{M}_{n}^{\prime})=|\lambda_{1}(\bm{M}_{n}^{\prime})|$ for sufficiently large $n$, meaning only the spectral behavior near the right edge $E_{+}$ is necessary to consider. A similar argument applies when $|E_{+}|<|E_{+}|$. While this relaxation may appear minor, the following example demonstrates its merit in extending the applicability of our theorem, even in the simplest case.
Example 1.
Consider the case $\alpha=1$ and $\bm{\Sigma}=\bm{I}_{p}$, $\bm{R}=-\bm{I}_{p}$. We can show that $\bm{M}_{n}^{\prime}=\phi^{-1/2}(\hat{\bm{\Sigma}}-\bm{I}_{p})$ satisfies Assumption 4 for all $n$ and $p$, but fails to satisfy Assumption 4 if $p>n$.
This demonstrates the relevance of introducing Assumption 4 when focusing on $\sigma(\bm{M}_{n}^{\prime})$. Specifically, in the case where $p>n$, the smallest eigenvalue, $\lambda_{p}(\bm{M}_{n}^{\prime})$, becomes isolated at the left edge $E_{-}=-\phi^{-1/2}$ and does not satisfy Assumption 4. However, we can show that when $p>n$, $|E_{+}|>|E_{-}|$, allowing Assumption 4 to hold consistently. Accordingly, we adopt Assumption 4 in our analysis of extreme singular values. Recalling that $T$ is the largest singular value of $\bm{\Sigma}-\bm{\Sigma}_{0}$, we are led to define the concept of an admissible pair.
@@
Theorem 3.4 (Anti-concentration inequality).
Under Assumption 1, for any admissible pair $(\bm{\Sigma},\bm{\Sigma}_{0})$ and $\delta\leq n^{-2/3}$, we have
$a(T,\delta)\prec n^{\frac{2}{3}}\delta.$ (36)
Equipped with these results, we present the universal consistency theorem for $T$. Recall the test in (5), where we reject $H_{0}$ if $T\geq\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{0}}(\alpha)$. Here $\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{0}}(\alpha)=\hat{q}^{\text{ub},B}_{\bm{\Sigma}_{0},\bm{\Sigma}_{0}}(\alpha)$ and $\hat{q}^{\text{ub},B}_{\bm{\Sigma},\bm{\Sigma}_{0}}(\alpha)$ denotes the empirical upper $\alpha$-th quantile of $T^{\text{ub},b}$, $b=1,\cdots,B$. To establish the theoretical validity of this universal bootstrap procedure, we provide a bound on the uniform Gaussian approximation error
@@
where $\mathbf{B}_{\text{op}}(\bm{\Sigma}_{0},t)$ is the operator norm ball centered at $\bm{\Sigma}_{0}$ with radius $t$. The probability in (37) is defined with respect to $\bm{X}$ and $\bm{Y}$, whose rows have covariance matrix $\bm{\Sigma}$, which may differ from $\bm{\Sigma}_{0}$. We summarize the results for universal bootstrap consistency as follows.
@@
Under Assumption 1, for any admissible pair $(\bm{\Sigma},\bm{\Sigma}_{0})$, we have the uniform Gaussian approximation bound
$\rho_{n}(\bm{\Sigma}_{0})\leq Cn^{-\delta}\to 0\quad\text{as}\quad n\to\infty,$ (38)
for some constants $C$, $\delta>0$.

### v2

@@ 3.1 Universality properties
Informally, Assumption (8) ensures that the extreme eigenvalues of $\bm{\Sigma}$ do not spread near the endpoints $E_{\pm}$. Similar assumptions have been made in the literature for the universality of $\hat{\bm{\Sigma}}$ (Bao et al., 2015; Knowles & Yin, 2017), with the Stieltjes transform and $\lambda_{i}(\bm{\Sigma})$ replaced by $\tilde{m}$ and $\lambda_{i}(\bm{\Sigma}(E_{\pm}))$, respectively.
@@
We denote by $\mathbb{P}^{\text{Gau}}$ and $\mathbb{E}^{\text{Gau}}$ the probability and expectation under which $\{Z_{ij}\}_{1\leq i\leq n,1\leq j\leq p}$ are independent standard Gaussian variables. With this, we present the following result.
@@
$\mathbb{P}(n^{\frac{2}{3}}($ $\lambda_{1}(\phi^{-1/2}\bm{M}_{n})-E_{+})\leq t-n^{-\epsilon})-n^{-\frac{1}{6}+C\epsilon}$
$\leq$ $\mathbb{P}^{\text{Gau}}(n^{\frac{2}{3}}(\lambda_{1}(\phi^{-1/2}\bm{M}_{n})-E_{+})\leq t)$ (9)
$\leq\mathbb{P}(n^{\frac{2}{3}}(\lambda_{1}(\phi^{-1/2}\bm{M}_{n})-E_{+})\leq t+n^{-\epsilon})+n^{-\frac{1}{6}+C\epsilon}.$
When $\gamma=1$, similar results also hold for $\lambda_{p}(\phi^{-1/2}\bm{M}_{n})$.
This universality theorem shows that the asymptotic distributions of $\lambda_{1}(\bm{M}_{n})$, $\lambda_{p}(\bm{M}_{n})$, and consequently $\sigma_{1}(\bm{M}_{n})=\|\bm{M}_{n}\|_{\text{op}}$, rely solely on the first two moments of $\bm{X}$. In contrast, Qiu et al. (2023) and Chernozhuokov et al. (2022) show that other widely-used norms of $\bm{M}_{n}$, such as $\|\bm{M}_{n}\|_{\text{F}}$ and $\|\bm{M}_{n}\|_{\text{sup}}$, are influenced by the first four moments of $\bm{X}$. This universal characteristic of the operator norm offers a straightforward yet effective framework for bootstrap. Before applying universality to the universal bootstrap, we give the following corollary of Theorem 3.1, which is of notable independent interest.
@@
Consider the case $\gamma>1$, $\bm{\Sigma}=\bm{I}_{p}$, $\bm{R}=\bm{0}$. Define $\mu=p$, $\sigma=p^{1/2}n^{-1/6}$. Under Assumption 1 for $\bm{Z}=(\bm{Z}_{1},\cdots,\bm{Z}_{n})^{T}$, we have as $n\to\infty$,
@@
where $\Rightarrow$” represents weak convergence, and $\text{TW}_{1}$ is the Tracy-Widom law of type $1$.
One of the central problems in random matrix theory is establishing the asymptotic distribution of the largest eigenvalue of a sample covariance matrix. In this context, (10) was derived for Gaussian $\bm{Z}$ in the setting where $p/n\to\infty$ (Karoui, 2003). However, for general distributions of $\bm{Z}$ where $\gamma>1$, no further results have been established. Our universality Theorem 3.1 addresses this significant gap in random matrix theory.
3.2 Consistency of universal bootstrap
We shall now provide the consistency results for our universal bootstrap using universality results established in Section 3.1. We remark that the theorems derived thus far apply specifically to the extreme eigenvalues $\lambda_{1}(\bm{M}_{n})$ and $\lambda_{p}(\bm{M}_{n})$. As we will show, extending these results to extreme singular values $\sigma_{1}(\bm{M}_{n})$ requires only weaker assumptions.
@@ 3.2 Consistency of universal bootstrap
When $\gamma=1,$ we require that there exists $\tau>0$ such that
@@
Assumption 4 is weaker than Assumption 4. While this relaxation appears minor, consider the case $\gamma=1$ and $\bm{\Sigma}=-\bm{R}=\bm{I}_{p}$. We can show that $\bm{M}_{n}^{\prime}$ satisfies Assumption 4 for all $n$ and $p$, but fails to satisfy Assumption 4 if $p>n$. This demonstrates the relevance of introducing Assumption 4 when analyzing extreme singular values. Recalling that $T$ is the largest singular value of $\bm{\Sigma}-\bm{\Sigma}_{0}$, we are led to define the concept of an admissible pair.
@@
Equipped with these results, we present the universal consistency theorem for $T$. To establish the theoretical validity of this universal bootstrap procedure, we provide a bound on the uniform Gaussian approximation error
@@
where $\mathbf{B}_{\text{op}}(\bm{\Sigma}_{0},t)=\big{\{}\bm{N}\ :\|\bm{N}-\bm{\Sigma}_{0}\|_{\text{op}}\leq t\big{\}}$ is the operator norm ball centered at $\bm{\Sigma}_{0}$ with radius $t$. The probability in (11) is defined with respect to $\bm{X}$ and $\bm{Y}$, whose rows have covariance matrix $\bm{\Sigma}$, which may differ from $\bm{\Sigma}_{0}$.
@@
Under Assumption 1, for any admissible pair $(\bm{\Sigma},\bm{\Sigma}_{0})$, we have the uniform Gaussian approximation bound for some constant $\delta>0$,
$\rho_{n}(\bm{\Sigma}_{0})\lesssim n^{-\delta}\to 0\quad\text{as}\quad n\to\infty.$ (12)